A buffer solution is composed of 0.1 mol NH3 and 0.1 mol NH4Cl in 1L. This...

  1. A buffer solution is composed of 0.1 mol NH3 and 0.1 mol NH4Cl in 1L. This solution is diluted 10 times. What is the pH change?
Answer #1

[NH37 = 0.1 mol/l = 0.1(M) [NHAC1] = 0 of mol/l = 0.1(M) .. The pol of the buffer soh can be given as - pOH = pkb (NH3) + log